OpenStudy (anonymous):

twelve less than one-fifth of a number is -7

TheSmartOne (thesmartone):

welcome to OS!! \[\frac{ 1 }{5}n-12=-7\] Can you solve from there?

TheSmartOne (thesmartone):

less than means subtract is means equal sign of means multiplying...

OpenStudy (anonymous):

1/5x-12=-7 5(1/5x)-5(12)=5(-7) x-60=-35 x=-35+60 x=25

TheSmartOne (thesmartone):

or you can add the 12 to both sides and you will get \[\frac{ 1 }{ 5 }x=5\] and then when you multiply 5 on both sides you will get \[x=25\]

TheSmartOne (thesmartone):

@alangonsalez